OpenStudy (anonymous):

This exercise is in a section about using trig substitution to integrate. I can use trig sub, but I'm just not seeing what initial substitution to use so I can apply trig sub. Can anyone give me a hint? I don't need (don't want) it solve to the end. At least not yet.

OpenStudy (anonymous):

See attached

OpenStudy (anonymous):

see this http://en.wikipedia.org/wiki/Trigonometric_substitution

OpenStudy (anonymous):

The trick is, how do we get this integral into the form\[dx \over \sqrt{u^2 +a^2}\]

OpenStudy (anonymous):

sorry du on top

OpenStudy (anonymous):

note, 16+x^6= (4)^2+(x^3)^2. or (x^3)^2+(4)^2.

OpenStudy (anonymous):

what will be the a ? and what will be the u?

OpenStudy (anonymous):

then, get the derivative of u.

OpenStudy (anonymous):

Thanks, that is just what I needed.

OpenStudy (anonymous):

most welcome:))